1)

According to Newtons Law of Cooling, if a body with temperature Ty is placed in surroundings with temperature To different fDON'T FORGET TO ROUND TO THE NEAREST INTEGER.
Answer #1

ang T (t) = To + +(T-T)ekt here T 0 Sure sreien unding tempo T, = Temp. of object here T, = ug tom? room tem fo - 68 from the10 - Afles minutes T(t) = 55 so , substitut and t=10 in eg. D lok 55= 68-ige Alge 55-68 lok 13 19 ree 13 -lok e. 19 fake bog-Ka 1 log ( 13 To 19 Put the value oblik) in eo 1 to bg () T(€) = 68-198 take t= 15 min. log (96) Tt) = 68-1970 19 68-19e le